OpenStudy (anonymous):

Carla takes 60 minutes to pick the mangoes from her backyard tree. Her brother can do it in 40 minutes. How long will it take them if they pick together?

hartnn (hartnn):

the general formula is (product)/(sum) so, whats 60*40/(60+40) =... ?

OpenStudy (hba):

\[\frac{ product }{ \sum },This \ is \ the formula \ \]

OpenStudy (hba):

So, \[\frac{ 60 \times 40 }{ 60+40 }=..\]